OpenStudy (anonymous):

What is the d value of the following arithmetic sequence? 27, 41, 55, 69, 83, 97….

OpenStudy (helder_edwin):

if it is an arithmetic sequence, then what happens with the difference of two consecutive elements? check that out!

OpenStudy (anonymous):

14

OpenStudy (anonymous):

I answered this one on previous post by another guy

OpenStudy (anonymous):

n=n-14 in descending order or n=n+14 in ascending order

OpenStudy (anonymous):

thank yu guys